Which aqueous solution is expected to have the highest boiling point?

  1. 1.5 FeCl2
  2. 3 CH3OH
  3.  2.5M C6H12O6
  4. 2.5M  NaCl

Grade:

2 Answers

Rituraj Tiwari
3 years ago
2.5mNaCl aqueous solution is expected to have the highest boiling point.
The elevation in the boiling point is directly proportional to the product
i×mwhere 'i' is the vant Hoff factor and 'm' is the molality of the solution.
The producti×mis maximum for2.5mNaCl.
For2.5mNaCli×m=2×2.5=5(since 1 molecule of NaCl on complete ionization gives 2 ions, the vant Hoff factor i is 2 for NaCl.)
ankit singh
3 years ago
2.5mNaCl aqueous solution is expected to have the highest boiling point.
The elevation in the boiling point is directly proportional to the product
i×mwhere 'i' is the vant Hoff factor and 'm' is the molality of the solution.
The producti×mis maximum for2.5mNaCl.
For2.5mNaCli×m=2×2.5=5(since 1 molecule of NaCl on complete ionization gives 2 ions, the vant Hoff factor i is 2